Peas heterozygous for three independently assorting genes were intercrossed. (example cross: Aa Bb Ccx Aa Bb Cc) What proportion of the offspring will be homozygous for all three recessive alleles? What proportion of the offspring will be homozygous for one gene and heterozygous for the other two?
Answer #1

a)

¼ x ¼ x ¼ =1/64

b)

i. 37/64: aa occurs in ¼ of all cases (64 x 0.25) = 16, 1. rest 64-16= 48, ¼ of the rest will be bb (48 x 0.25) = 12. 2. Rest 48-12= 36, ¼ of the rest will be cc (36 x 0.25) = 9. 3. At least is additive 16 +12 +9 =37/64.